OpenStudy (roxanne):

for each equation find three ordered pair solutions by completing the table. then use ordered pairs to graph the equation y=1/3x x is 0 x is 6 x is -3

OpenStudy (anonymous):

k whats the question?

OpenStudy (anonymous):

substitute those numbers in for x the answer will be the y value

OpenStudy (anonymous):

then graph

OpenStudy (anonymous):

y=1/3(0) = 0

OpenStudy (anonymous):

so (0,0)

OpenStudy (anonymous):

y=1/3(6)=2 (6,2)

OpenStudy (anonymous):

y=1/3(-3)=-1 (-3,-1)

OpenStudy (anonymous):

now just graph the points